OpenStudy (oayala):

anybody know the steps to solving this? , Jake sold 39 tickets to the school fair and Jeanne sold 12 tickets. What is the ratio, in simplest form, of the number of tickets Jeanne sold to the number of tickets Jake sold? 4/13 12/39 39/12 13/4

OpenStudy (welshfella):

12/39 must be changed to its simplest form Hint Factor the 2 numbers eg 12 = 3*4 39 = ?

OpenStudy (oayala):

3x13 equals 39

jhonyy9 (jhonyy9):

yes sure 39=3*13 so than you rewrite this rationality 12/39 = ?

jhonyy9 (jhonyy9):

oayala do you understand it ? you ned rewrite this numbers of fraction for you can simplify it and so in this way will get the right answer sure

jhonyy9 (jhonyy9):

you need rewrite it sorry

jhonyy9 (jhonyy9):

12/39 = (3*4) / ?

OpenStudy (jane---):

4/13

OpenStudy (goginal):

the answer is 4/13 because the ratio is 12/39 and you find a number that goes into both 12 and 39 so you can simplify 12/39 and the number 3 can go into 12 and 39 so you divide 39 by 3 which is 13 and you also divide 12 by 3 which is 4 so your answer is 12/39 and 4/13 in simplest form